OpenStudy (anonymous):

dealing with logarithmic functions ln1/2=-0.693....... write the log. equation in exponential form dealing with logarithmic functions ln1/2=-0.693....... write the log. equation in exponential form @Mathematics

OpenStudy (asnaseer):

remember that if:\[\ln{x}=y\]then this implies:\[e^y=x\]

OpenStudy (asnaseer):

because:\[\ln{x}=\log_e{x}\]

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

yep thanks

OpenStudy (asnaseer):

yw